Dry ice is carbon dioxide in the solid state .0300 moles of dry ice is placed in a 5.00 L chamber that is maintained at 35.1oC. What is the pressure in the chamber after all of the dry ice has sublimed?
1
Expert's answer
2020-05-28T10:33:11-0400

According to the ideal gas equation:

p × V = n × R × T

where p - pressure, V - volume, n - number of moles, R - gas constant (8.02578 × 10-2 L atm / K mol), T - temperature (35.1 °C = 308.25 K).

From here:

p = n × R × T / V = 0.0300 mol × 8.02578 × 10-2 L atm / K mol × 308.25 K / 5.00 L = 0.148 atm


Answer: 0.148 atm